OpenStudy (anonymous):

The community 10K run/walk event awards ribbons to the top four finishers in each age category. If 20 people in the 15-19 age range register for the race, how many ways are there to arrange the top four finishers?

OpenStudy (anonymous):

@Brostep0s

OpenStudy (anonymous):

20C4=20*19*18*17/2*3*4

OpenStudy (anonymous):

What about the answer bit?

OpenStudy (anonymous):

20C4=20*19*18*17/2*3*4=4845 possible arrangements.